Circular No. 5674

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CYGNI 1992 M. A. Greenhouse, National Air and Space Museum (NASM); C. E.

Woodward, University of Wyoming; J. Fischer, Naval Research Laboratory; and H. A. Smith, NASM, communicate: “Spectroscopy of N Cyg 1992 obtained on Oct. 18.35 UT (243 days after outburst) with the NASA Infrared Telescope Facility (+ CGAS) reveals the following infrared coronal lines (wavelengths listed are the observed values, with uncertainty in last digit given in parentheses): [Al VI] 3.657(2)-μm, 2s^{2}2p^{4} ^{3}P_{1} −^{3}P_{0}; [Al VIII] 3.699(2)-μm, 2s^{2}2p^{2} ^{3}P_{1} −^{3}P_{2}; [Mg VIII] 3.030(2)-μm, 2s^{2}2p^{2} ^{2}P_{1/2} −^{3}P_{3/2}; and a feature at 3.09(1) μm tentatively identified as [Ca IX] 3s3p ^{3}P_{1} −^{3}P_{2}. Their integrated intensities were (in units of 10^{−10} erg s^{−1} cm^{−2}): [Al VI], 0.12 ± 0.01; [Al VIII], 0.020 ± 0.008; [Mg VIII], 5.20 ± 0.04; [Ca IX], 0.08 ± 0.02. The observed free-free flux density was 0.50 ± 0.09 Jy at 3.6 μm. The lines were velocity-resolved at ∼ 2600 km/s FWHM. The Al and Mg lines have been observed previously in two other novae (Greenhouse et al. 1990, Ap.J. 352, 307), and suggest that N Cyg 1992 occurred on an ONeMg white dwarf (Starrfield et al. 1986, Ap.J. 303, L5). The Ca transition has not been reported previously in an astronomical source. We note that the infrared Ca transition discussed here should be accompanied by a 3s^{2} ^{1}S_{0} − 3s3p ^{3}P_{2} magnetic quadrapole transition at 67.61 nm. The [Al VI] line was first observed in N Cyg 1992 on Sept. 9.29 (IAUC 5612) at an intensity of 1.10 × 10^{−12} erg s^{−1} cm^{−2}. The increased [Al VI] intensity and the large [Mg VIII]/[Al VI] intensity ratio reported here suggests that the coronal line phase is near maximum. The observed line width suggests that no appreciable deceleration of the ejecta has occurred since the onset of the optically thin emission-line phase (e.g., Pendelton et al., IAUC 5544). Additional infrared spectroscopy covering M, N, and Q bands are needed and should reveal similar emissions.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 This comet is again in outburst, as reported by I. Ferrin, National Ob-

servatory of Venezuela, and A. Nakamura, Kuma, Japan, who provide the following total visual magnitude estimates: Nov. 24.69 UT, [14.5 (Nakamura, 0.60-m reflector); 30.70, 13.3 (Nakamura); Dec. 2.72, 13.1 (Nakamura); 5.03, 14.1 (Ferrin, 1-m Schmidt telescope + CCD).

1992 December 10 Daniel W. E. Green